OpenStudy (anonymous):

the sum of the measrues of two angles is 180 degrees one angle measures 3a,and the other angle measures 2a-25find a.then find the measure of each angle.

OpenStudy (anonymous):

3a + 2a -25 =180 5a =180+25 5a = 205 a = 41 then put value of a that is 3a = 3*41= 123 2a-25= 57 so you gft your answer.